OpenStudy (anonymous):

9-5n+1=15 n= i still don't understand

OpenStudy (anonymous):

You need to solve for "n".

OpenStudy (anonymous):

We want to get n alone on one side of the equation. First combine like terms. 9 - 5n + 1 = 15 (9 + 1) 10 - 5n = 15

OpenStudy (anonymous):

Subtract the 10 from both sides. Remember to keep the symbol in front of each number/term. 10 - 5n = 15 -10 -10 -5n = 5

OpenStudy (anonymous):

Now divide both sides by -5. \[\frac{ -5n }{ -5 } = \frac{ 5 }{ -5 }\]n = -1

OpenStudy (anonymous):

Hopefully this helped you. (:

OpenStudy (anonymous):

thx.

OpenStudy (anonymous):

u think u can help with one more q. @jazy ?

OpenStudy (anonymous):

Sure!